OpenStudy (kobeni-chan):

Using a directrix of y = -2 and a focus of (2, 6), what quadratic function is created?

OpenStudy (michele_laino):

I think it is a parabola

OpenStudy (kobeni-chan):

 my answer choices are all in this format

OpenStudy (kobeni-chan):

and I know that it'll be - 2 at the end from the directrix

OpenStudy (kobeni-chan):

actually now that I'm looking at all the choices, the first fraction part has to be either 1/8 or -1/8, but I don't remember how to find that part

OpenStudy (michele_laino):

if I write your parabola as below: \[y=ax ^{2}+bx+c\] then coefficients a, b, c are related to the coordinates of focus and to the equation of directrix, by the subsequent formulas: \[-\frac{ b }{ 2a }=2\] \[\frac{ 1-\Delta }{ 4a }=6\] \[-\frac{ 1+\Delta }{ 4a }=-2\] where: \[\Delta =b ^{2}-4ac\]

OpenStudy (kobeni-chan):

how can I find the values for a, b, and c? I don't think I can put them in the form of y=ax^2 + bx + c

OpenStudy (michele_laino):

don't worry after we found a, b and c we will get the parabola in your form Now I'm finding a, b and c

OpenStudy (kobeni-chan):

ok

OpenStudy (michele_laino):

from the preceding formulas, we get: \[a=\frac{ 1 }{ 16 },b=-\frac{ 1 }{ 4 },c=\frac{ 9 }{ 4 }\] so our parabola is: \[y=\frac{ 1 }{ 16 }x ^{2}-\frac{ 1 }{ 4 }x+\frac{ 9 }{ 4 }=\frac{ 1 }{ 16 }(x ^{2}-4x+36)\]

OpenStudy (michele_laino):

\[=\frac{ 1 }{ 16 }(x ^{2}-4x+4+32)=\frac{ 1 }{ 16 }[(x-2)^{2}+32]=\]

OpenStudy (michele_laino):

\[\frac{ 1 }{ 16 }[(x-2)^{2}]+\frac{ 32 }{ 2 }=\frac{ 1 }{ 16 }[(x-2)]^{2}+2\]

OpenStudy (michele_laino):

so your parabola is: \[y=\frac{ 1 }{ 16 }[(x-2)]^{2}+2\]

OpenStudy (kobeni-chan):

ok thank you! :)

OpenStudy (michele_laino):

thank you! :)

OpenStudy (michele_laino):

I'm sorry, but I don't know a simpler method!

OpenStudy (kobeni-chan):

lol that's ok